# Difference-in-Differences Designs: A Practitioner’s Guide

### v1

@@ 4 Incorporating covariates into 2$\times$2 DiD
So far, we have focused on $2\times 2$ DiD designs that do not leverage any information about covariates, but researchers frequently incorporate them into DiD analyses in one of three ways: checking for balance in variables thought to influence $Y_{i,t}(0)$, controlling for those variables in the main estimates, and estimating treatment effect heterogeneity. For example, in the absence of Medicaid expansion, mortality rates likely would have evolved differently in poorer and richer counties and certainly did prior to 2014 (Currie2016). Therefore, parallel trends may not be plausible if poverty rates differ between expansion and non-expansion counties. If they do, then one may want to “control for” poverty rates when estimating $ATT$ parameters. Finally, because Medicaid expansion reached more people in higher-poverty counties, its average effects on overall mortality may be larger there. This heterogeneity may be of interest in its own right or may be used to assess the plausibility of the overall DiD design.
@@ 4.1 Covariate balance: Is unconditional parallel trends plausible?
Assumption PT is fundamentally untestable, as it contains an unobserved counterfactual component. Therefore, “tests” of these assumptions are necessarily indirect and rely on other observed variables thought to be related to untreated potential outcome trends. For example, during the 2010s, area-level demographics and economic conditions were strongly correlated with mortality levels and trends. If these relationships would have held in the absence of Medicaid expansion, and if expansion and non-expansion counties differ in those demographic and economic characteristics, then the parallel trends assumption (3.3) may fail to hold. Checking balance in observable determinants of changes in $Y_{i,t}(0)$ is thus a common and sensible way to evaluate parallel trends.
Most DiD analyses check for balance across groups in baseline covariate levels ($\mathop{}\!\mathbb{E}_{\omega}[X_{i,t=1}|D_{i}=1]-\mathop{}\!\mathbb{E}_{\omega}[X_{i,t=1}|D_{i}=0]$) or covariate trends before and after treatment ($\mathop{}\!\mathbb{E}_{\omega}[\Delta X_{i,t=2}|D_{i}=1]-\mathop{}\!\mathbb{E}_{\omega}[\Delta X_{i,t=2}|D_{i}=0]$, where $\Delta X_{i,t=2}=X_{i,t=2}-X_{i,t=1}$). We consider the following variables, $X_{i,t}$: the percentages of a county’s population that are female, white, or Hispanic; the unemployment rate; the poverty rate; and county-level median income (in thousands of dollars). [^11] We focus on these variables for convenience. Borgschulte2020 use a LASSO procedure that selects more and different covariates to include in their analysis. We replicate their findings when we follow their methodology but diverge here for the sake of brevity. Panel A of Table 4 reports averages of these variables by group in 2013 with and without population weights. We also report a measure of imbalance that is comparable across variables: the normalized difference in means between treatment and comparison group (imbensrubin2015, Chapter 14),
@@
where $\overline{X}_{\omega,T}$ and $\overline{X}_{\omega,C}$ are the sample weighted or unweighted averages for the treatment and comparison groups, respectively, and $S_{\omega,T}^{2}$ and $S_{\omega,C}^{2}$ are the sample weighted or unweighted variances of the covariates for the treatment and comparison group. As a general rule of thumb, values of the normalized difference in excess of 0.25 in absolute value indicate a potentially problematic imbalance between the two groups (imbensrubin2015, page 277).
@@
  • Notes: This table reports the covariate balance between adopting and non-adopting states. In the top panel, we report the averages and standardized differences of each variable, measured in 2013, by adoption status. All variables are measured in percentage values, except for median household income, which is measured in thousands of U.S. dollars. In the bottom panel we report the average and standardized differences of the county-level long differences between 2014 and 2013 of each variable. We report both weighted and unweighted measures of the averages to correspond to the different estimation methods of including covariates in a 2 $\times$ 2 setting.
@@
We find meaningful imbalance in several baseline measures. Expansion counties in 2013 were whiter and had a higher unemployment rate despite lower poverty and higher median income than non-expansion counties. Because DiD uses changes in outcomes, sometimes researchers argue that the effect of pre-treatment variables is differenced out. This logic does not hold, though, if baseline covariates are related to untreated potential outcome trends themselves. The imbalance in panel A of Table 4 will lead to violations of parallel trends to the extent that counties starting out with different racial composition or income distributions would have had different mortality trends even absent Medicaid expansion.
@@
Nevertheless, balance in covariate changes can be informative about parallel trends as well. Panel B of Table 4 reports average changes by group between 2013 and 2014 as well as normalized differences. Many of the imbalances evident in baseline levels change, or even flip signs, when measured in changes. Unemployment, for example, was higher in expansion states in 2013 but fell faster. To the extent that these changes are important determinants of $\Delta Y_{i,t}(0)$, then these results could suggest that Assumption PT is violated.
@@
Why do we say “could”? A major challenge in interpreting cross-group gaps in $\Delta X_{i,t}$ involves deciding which variables are truly covariates and which are mechanisms/outcomes. If an element of $X_{i,t}$ cannot be affected by the treatment, it is a (strictly exogenous) covariate, and differential changes in exogenous covariates may indicate a PT violation. Since the treatment cannot have caused $X_{i,t}$ to change (by assumption), something else that differs across groups and over time must have. Since little research suggests an effect of Medicaid expansion on unemployment, this may be a good assumption. On the other hand, if Medicaid expansion can change the demographic and economic composition of its counties, then differential changes in these variables may actually be a consequence of the expansion itself. [^12] In fact, comparing mean covariate changes in expansion and non-expansion is the same as using $X_{i,t}$ as the outcome in a $2\times 2$ DiD estimator. If so, then differential post-treatment changes in them would not necessarily indicate a parallel trends violation; they could partially reflect a causal effect. Like the plausibility of Assumption PT itself, whether something is a covariate or a mechanism is not a data question per se. It requires context-specific knowledge (or assumptions) about how treatment works.
@@ 4.2 DiD with covariates: Identification under conditional parallel trends
Let $X_{i}$ be a vector of observed determinants of changes in $Y_{i,t}(0)$. Here, we purposefully omit the time subscript on $X_{i}$ because the covariates in this section can be time-invariant, such as fixed variations or baseline values ($X_{i,t=1}$), or time-varying in the sense of including values from in the second period, $X_{i,t=2}$. The empirical content of a “new” identification assumption that incorporates $X_{i}$, henceforth conditional parallel trends (CPT) assumption, is formalized as follows.
@@
Assumption CPT has the same structure as Assumption PT but states that PT holds within each covariate-specific stratum rather than across the whole population. With respect to the baseline covariates in Table 4, this assumption amounts to assuming parallel trends in $Y_{i,t}(0)$ between expansion and non-expansion counties with the same female, white, and Hispanic shares as well as unemployment and poverty rates, and median income in 2013. This does not restrict how $Y_{i,t}(0)$ changes in different covariate strata, and it is not necessary to specify or estimate whether mortality rates would have gone up or down for counties with specific characteristics. It only imposes that these covariate-specific trends are common between expansion and non-expansion counties.
@@
The conditional (weighted) probability of belonging to the treatment group, given observed covariates $X_{i}$ that are determinants of untreated potential outcome growth, is uniformly bounded away from zero and one. That is, for some $\epsilon>0$, $\epsilon<P_{\omega}[D_{i}=1|X_{i}]<1-\epsilon$.
@@
$=\mathop{}\!\mathbb{E}_{\omega}[\Delta Y_{i,t=2}|D_{i}=1\Big{]}-\mathop{}\!\mathbb{E}_{\omega}\Big{[}\mathop{}\!\mathbb{E}_{\omega}[\Delta Y_{i,t=2}|X_{i},D_{i}=0]\Big{|}D_{i}=1\Big{]}$ (4.2)
The first line restates the definition of $ATT(2)$, and the second line uses the law of iterated expectations to write the counterfactual mean for the whole treatment group as an average of counterfactual means conditional on $X_{i}$. These quantities are exactly the ones that appear in the conditional parallel trends assumption (and the overlap condition). Therefore, just like in section 3, the third line uses Assumption CPT to rewrite the conditional counterfactuals in terms of observable population quantities (under Assumption NA). Equation (4.2) then uses the law of iterated expectations again to group terms and express the $ATT(2)$ in terms of observed variables $(Y_{i,t=2},Y_{i,t=1},G_{i},X_{i})$, i.e., it establishes that $ATT(2)$ is nonparametrically identified under our assumptions. This expression has a clear intuition: the $ATT(2)$ is equal to the path of outcomes experienced by the treated group (the term on the left) minus the average path of outcomes in the comparison group for each value of the covariates, averaged over the treated group’s distribution of covariates (the term on the right).
@@ 4.3 DiD estimation with covariates: TWFE
Unlike in an unconditional DiD, moving from the population identification result in equation (4.2) to sample analogs is a challenge unless the covariates are discrete and the conditional expectations themselves are easily calculable. With continuous covariates, or many discrete ones, it may not be feasible to construct $\mathop{}\!\mathbb{E}_{\omega}[\Delta Y_{i,t=2}|X_{i},D_{i}=0]$. Conditional DiD estimation, therefore, uses additional econometric techniques to bridge this gap. We begin, however, by discussing how regression DiD estimators that include covariates relate to the assumptions used for identification in (4.2).
Because the TWFE specification in (3.7) recovers the $ATT(2)$ in $2\times 2$ DiD setups without covariates, it is natural to extend this logic to regressions with covariates. Indeed, this is by far the most popular approach adopted by practitioners arguably because it is both easy and familiar. A typical regression specification is
$Y_{i,t}=\theta_{t}+\eta_{i}+\beta_{treat}D_{i,t}+X_{i,t}^{\prime}\beta_{covs}+e_{i,t},$ (4.3)
where the unit and time fixed effects, treatment status, and covariates have already been defined, $e_{i,t}$ is an error term, and $\beta_{treat}$ is interpreted as the parameter of interest. A related specification explicitly controls for baseline covariates by replacing $X_{i,t}$ with interactions of the pre-treatment covariates and a post-treatment dummy,
@@
In Table 5, we report the OLS and weighted least squares estimates of the unconditional $2\times 2$ DiD estimate, $\beta_{covs}$ from (4.4), $\beta_{covs,2}$, from (4.3), and their cluster-robust standard errors using the covariates from Table 4.
@@
Unweighted Weighted No Covs $X_{i,t=2013}$ $X_{i,t}$ No Covs $X_{i,t=2013}$ $X_{i,t}$ (1) (2) (3) (4) (5) (6) Medicaid Expansion 0.12 -2.35 -0.49 -2.56 -2.56 -1.37 (3.75) (4.29) (3.83) (1.49) (1.78) (1.62) Notes: This table reports the regression 2 $\times$ 2 DiD estimate comparing counties that expand Medicaid in 2014 to counties that do not expand Medicaid, adjusting for the inclusion of covariates (percent female, percent white, percent hispanic, the unemployment rate, the poverty rate, and median household income). Columns 1-3 report unweighted regression results, while columns 4-6 weight by county population aged 20-64 in 2013. Columns 1 and 4 report results for expansion states without covariates, columns 2 and 5 adjust for the baseline levels of the covariates in 2013, and columns 3 and 6 control for the time-varying covariate values in 2014 and 2013. Standard errors (in parentheses) are clustered at the county level.
@@
Although only one covariate-adjusted estimate in Table 5 is (marginally) statistically significant, the point estimates differ noticeably. In the unweighted case, adjusting for the 2013 levels of the covariates decreases the estimated effect of Medicaid expansion on short-run mortality rates from a point estimate of roughly 0.12 to -2.35. However, if we include their time-varying values instead, we estimate an effect of -0.49, a large difference. We find a similar result when using weighted regressions; while the coefficient remains fairly constant (-2.56) when using 2013 values of the covariates, it attenuates to -1.37 if we use (4.3).
@@
The first thing that is clear from these representations is that because time-invariant variables drop out of equation (4.3), a TWFE specification can only account for differential trends related to baseline covariate levels if they enter as interactions with the post-treatment dummy as in equation (4.4). The exact regression specification, therefore, determines the implied conditional parallel trends assumption. Controlling for annual poverty rates really means controlling for poverty changes, and areas that are poor are not the same as areas that are becoming poor.
@@
Suppose we have decided on which variables to include in a conditional parallel trends assumption and whether to measure them in levels or changes. If Assumptions CPT and SO hold with respect to this set of covariates, does $\beta_{treat}$ recover the $ATT(2)$? In the DiD context, Caetano2024 tackles exactly this question. They show that $\beta_{treat}$ equals a weighted average of conditional average treatment effects, defined as $ATT_{x_{k}}(2)\equiv\mathop{}\!\mathbb{E}_{\omega}\big{[}Y_{i,t=2}(1)-Y_{i,t=2}(0)|D_{i}=1,X_{i}=x_{k}\big{]}$, with weights that may not be convex, plus three bias terms reflecting misspecification either in the set of control variables or the fact that they are included linearly. These conclusions relate to recent findings about the properties of regression estimators in the presence of heterogeneity in other contexts, including instrumental variables (Mogstad_Torgovitsky_2024), cross-sectional designs (Angrist_1998_ECMA; Aronow_Samii_2015; Sloczyski2022; Goldsmith-Pinkham_Hull_Kolesar_2024_AER), and panel data (Goodman2021; deChaisemartin2020; Sun2021; Poirier_Sloczyski_2024).
The weighting results suggest that even when the covariates are correctly selected, measured, and added with the correct functional form, $\beta_{treat}$ could be negative even when $ATT_{X_{i}}(2)$ is positive for all values of covariates. Short of this extreme sign-reversal case, $\beta_{treat}$’s weighting scheme generally does not yield the $ATT(2)$ target parameter and instead puts too much weight on $ATT_{X_{i}}(2)$ for $X_{i}$’s that are relatively uncommon among the treated group relative to the untreated group and puts too little weight on $ATT_{X_{i}}(2)$ for $X_{i}$’s that are relatively common among the treated group relative to the untreated group (Sloczyski2022, Caetano2024).
@@
In other words, (4.3) implicitly rules out that treatment effects can vary across covariate-strata, which makes the weighting issues identified by Caetano2024 irrelevant to the interpretation of $\beta_{treat}$. Research on the Medicaid expansion using data on mortality rates by income shows clear evidence of heterogeneous effects (miller2021medicaid; wyse2024), making this implicit restriction implausible.
@@
As discussed in the introduction, we believe forward-engineering with DiD provides clear benefits compared to the more standard reverse-engineering approach that fixes the econometric tool and a regression specification (e.g., TWFE) and tries to back out a reasonable (weakly) causal interpretation of the regression coefficient in the presence of unmodelled treatment effect heterogeneity that leads to model misspecification. When using the forward-engineering approach, changes in the estimation procedures lead to changes in estimates, but we know exactly what is being estimated and under which assumptions. In this sense, and contrary to the reverse-engineering approach, it does not move the goalpost, leading to a more transparent analysis, with less ambiguity about the identification assumptions and the interpretations of the parameters. A drawback is that one may need to learn and use new econometric techniques, though we view this as a minor step compared to its potential benefits.
We now turn to estimators that take this forward-engineering approach, which we will also adapt to DiD designs that are more complex than the $2\times 2$ setup heretofore discussed.

### v2

@@ 4 Incorporating covariates into 2$\times$2 DiD
So far, we have focused on $2\times 2$ DiD designs that do not leverage any information about covariates, but researchers frequently incorporate them into DiD analyses in one of three ways: checking for balance in variables thought to influence $Y_{i,t}(0)$, controlling for those variables in the main estimates, and estimating treatment effect heterogeneity. For example, in the absence of Medicaid expansion, mortality rates likely would have evolved differently in poorer and richer counties; they certainly did before 2014 (Currie2016). Therefore, parallel trends may fail if poverty rates differ between expansion and non-expansion counties. If they do, then one may want to “control for” poverty rates when estimating $ATT$ parameters. Finally, because Medicaid expansion reached more people in higher-poverty counties, its average effects on overall mortality may be larger there. This heterogeneity may be of interest in its own right or may be used to assess the plausibility of the overall DiD design.
@@ 4.1 Covariate balance: Is unconditional parallel trends plausible?
Assumption PT is fundamentally untestable, as it contains an unobserved counterfactual component. Therefore, “tests” of these assumptions are necessarily indirect and rely on other observed variables thought to be related to untreated potential outcome trends. For example, during the 2010s, demographics and economic conditions were strongly correlated with mortality levels and trends. If these relationships would have held in the absence of Medicaid expansion, and if expansion and non-expansion counties differ in those demographic and economic characteristics, then the parallel trends assumption (3.3) may fail to hold. Checking balance in observable determinants of changes in $Y_{i,t}(0)$ is thus a common and sensible way to evaluate parallel trends.
Most DiD analyses check for balance across groups in baseline covariate levels ($\mathop{}\!\mathbb{E}_{\omega}[X_{i,t=1}|D_{i}=1]-\mathop{}\!\mathbb{E}_{\omega}[X_{i,t=1}|D_{i}=0]$) or covariate trends before and after treatment ($\mathop{}\!\mathbb{E}_{\omega}[\Delta X_{i,t=2}|D_{i}=1]-\mathop{}\!\mathbb{E}_{\omega}[\Delta X_{i,t=2}|D_{i}=0]$, where $\Delta X_{i,t=2}=X_{i,t=2}-X_{i,t=1}$). We consider the following variables, $X_{i,t}$: the percentages of a county’s population that are female, white, or Hispanic; the unemployment rate; the poverty rate; and county-level median income (in thousands of dollars). [^10] We focus on these variables for convenience. Borgschulte2020 use a LASSO procedure that selects more and different covariates to include in their analysis. We replicate their findings when we follow their methodology but diverge here for the sake of brevity. The top panel of Table 4 reports averages of these variables by group in 2013 with and without population weights. We also report a measure of imbalance that is comparable across variables: the normalized difference in means between the treatment and comparison groups (imbensrubin2015, Chapter 14),
@@
where $\overline{X}_{\omega,T}$ and $\overline{X}_{\omega,C}$ are the sample weighted or unweighted averages for the treatment and comparison groups, respectively, and $S_{\omega,T}^{2}$ and $S_{\omega,C}^{2}$ are the sample weighted or unweighted variances of the covariates for the treatment and comparison group. As a general rule of thumb, values of the normalized difference in excess of 0.25 in absolute value indicate a potentially problematic imbalance between the two groups (imbensrubin2015, page 277). [^11] As discussed in Austin2009_Balance, the normalized difference was initially proposed in the psychological literature and is sometimes referred to as Cohen’s effect size index. In this context, normalized differences of 0.2, 0.5, and 0.8 are sometimes used to represent small, medium, and large imbalances; however, normalized differences as small as 0.1 are sometimes considered worrisome, depending on how important the covariate in question is. Ultimately, there is no universally accepted threshold for what value indicates important imbalances. See Ho_Imai_King_Stuart_2007_matching and Austin2009_Balance for additional discussions.
@@
  • This table reports the covariate balance between 978 counties in states that expanded Medicaid in 2014 and 1,222 counties in states that did not expand by 2019. In the top panel, we report the averages and standardized differences of each variable, measured in 2013, by adoption status. All variables are measured in percentage values, except for median household income, which is measured in thousands of U.S. dollars. In the bottom panel we report the average and standardized differences of the county-level long differences between 2014 and 2013 of each variable. We report both weighted and unweighted measures of the averages to correspond to the different estimation methods of including covariates in a 2 $\times$ 2 setting.
@@
We find meaningful imbalance in several baseline measures. Expansion counties in 2013 were whiter and had higher unemployment rates despite lower poverty and higher median income than non-expansion counties. Because DiD uses changes in outcomes, researchers sometimes argue that the effect of pre-treatment variables is differenced out. This logic does not hold, though, if baseline covariates are related to untreated potential outcome trends themselves. The imbalance in the top panel of Table 4 will lead to violations of parallel trends to the extent that counties with different racial composition or income distributions would have had different mortality changes even without Medicaid expansion.
@@
Nevertheless, checks of balance in covariate changes can be informative about parallel trends as well. The bottom panel of Table 4 reports average changes by group between 2013 and 2014 as well as normalized differences. Many of the imbalances evident in baseline levels change, or even flip signs, when measured in changes. Unemployment, for example, was higher in expansion states in 2013 but fell faster. To the extent that these changes are important determinants of $\Delta Y_{i,t}(0)$, then these results could suggest that Assumption PT is violated.
@@
Why do we say “could”? A major challenge in interpreting cross-group gaps in $\Delta X_{i,t}$ involves deciding which variables are truly covariates and which are mechanisms/outcomes. If an element of $X_{i,t}$ cannot be affected by the treatment, it is a (strictly exogenous) covariate, and differential changes in exogenous covariates may indicate a PT violation. Since the treatment cannot have caused $X_{i,t}$ to change (by assumption), something else that differs across groups and over time must have. Since little research suggests an effect of Medicaid expansion on unemployment, this may be a good assumption. On the other hand, if Medicaid expansion can change the demographic and economic composition of its counties, then differential changes in these variables may actually be a consequence of the expansion itself. [^12] In fact, comparing mean covariate changes in expansion and non-expansion is the same as using $X_{i,t}$ as the outcome in a $2\times 2$ DiD estimator. If so, then differential post-treatment changes in them would not necessarily indicate a parallel trends violation; they could partially reflect a causal effect. As with the plausibility of Assumption PT itself, whether something is a covariate or a mechanism is not a data question per se. It requires context-specific knowledge about how the treatment works.
@@ 4.2 DiD with covariates: Identification under conditional parallel trends
Let $X_{i}$ be a vector of observed determinants of changes in $Y_{i,t}(0)$. Here, we purposefully omit the time subscript on $X_{i}$ because the covariates in this section can be time-invariant, such as fixed variations or baseline values ($X_{i,t=1}$), or time-varying in the sense of including values from the second period, $X_{i,t=2}$. The empirical content of a “new” identification assumption that incorporates $X_{i}$, henceforth conditional parallel trends (CPT) assumption, is formalized as follows.
@@
Assumption CPT has the same structure as Assumption PT but states that PT holds within each covariate-specific stratum rather than across the whole population. With respect to the baseline covariates in Table 4, this amounts to assuming parallel trends in $Y_{i,t}(0)$ between expansion and non-expansion counties that in 2013 had the same female, white, and Hispanic shares, as well as the same unemployment and poverty rates and median income. This does not restrict how $Y_{i,t}(0)$ changes in different covariate strata, nor is not necessary to estimate these trends. Assumption CPT only requires that covariate-specific trends are common between expansion and non-expansion counties.
@@
The conditional (weighted) probability of belonging to the treatment group, given observed covariates $X_{i}$, which are determinants of untreated potential outcome growth, is uniformly bounded away from zero and one. That is, for some $\epsilon>0$, $\epsilon<P_{\omega}[D_{i}=1|X_{i}]<1-\epsilon$.
@@
$=\mathop{}\!\mathbb{E}_{\omega}[\Delta Y_{i,t=2}|D_{i}=1\Big{]}-\mathop{}\!\mathbb{E}_{\omega}\Big{[}\mathop{}\!\mathbb{E}_{\omega}[\Delta Y_{i,t=2}|X_{i},D_{i}=0]\Big{|}D_{i}=1\Big{]}.$ (4.2)
The first line restates the definition of $ATT(2)$, and the second line uses the law of iterated expectations to write the counterfactual mean for the whole treatment group as an average of counterfactual means conditional on $X_{i}$. These quantities are exactly the ones that appear in the conditional parallel trends assumption (and the overlap condition). Therefore, as in section 3, the third line uses Assumption CPT to rewrite the conditional counterfactuals in terms of observable population quantities under no anticipation (Assumption NA). Equation (4.2) then uses the law of iterated expectations again to group terms and express $ATT(2)$ in terms of observed variables $(Y_{i,t=2},Y_{i,t=1},G_{i},X_{i})$; that is, it establishes that $ATT(2)$ is nonparametrically identified under our assumptions. This expression has a clear intuition: the $ATT(2)$ is equal to the path of outcomes experienced by the treated group (the term on the left) minus the average path of outcomes in the comparison group for each value of the covariates, averaged over the treated group’s distribution of covariates (the term on the right).
@@ 4.3 DiD estimation with covariates: TWFE
Moving from the population identification result in equation (4.2) to sample analogs is a challenge unless the covariates are discrete and the conditional expectations themselves are easily calculable. This difficulty does not arise in an unconditonal DiD. With continuous covariates, or many discrete ones, it may not be feasible to construct $\mathop{}\!\mathbb{E}_{\omega}[\Delta Y_{i,t=2}|X_{i},D_{i}=0]$. Conditional DiD estimation, therefore, uses additional econometric techniques to bridge this gap. We begin, however, by discussing how regression DiD estimators that include covariates relate to the assumptions used for identification in (4.2).
Because the TWFE specification in (3.7) recovers the $ATT(2)$ in $2\times 2$ DiD setups without covariates, it is natural to extend this logic to regressions with covariates. Indeed, this is by far the most popular approach adopted by practitioners, arguably because it is both easy and familiar. A typical regression specification is
$Y_{i,t}=\theta_{t}+\eta_{i}+\beta_{treat}D_{i,t}+X_{i,t}^{\prime}\beta_{covs}+e_{i,t}.$ (4.3)
where the unit and time fixed effects, treatment status, and covariates have already been defined, $e_{i,t}$ is an error term, and $\beta_{treat}$ is interpreted as the parameter of interest. A related specification explicitly controls for baseline covariates by replacing $X_{i,t}$ with interactions of the pre-treatment covariates $X_{i,t=1}$ and a post-treatment dummy,
@@
In Table 5, we report the OLS and weighted least squares estimates of the unconditional $2\times 2$ DiD estimate, $\beta_{covs}$ from (4.4), $\beta_{covs,2}$ from (4.3), and their cluster-robust standard errors, using the covariates from Table 4.
@@
Unweighted Weighted No Covs $X_{i,t=2013}$ $X_{i,t}$ No Covs $X_{i,t=2013}$ $X_{i,t}$ (1) (2) (3) (4) (5) (6) Medicaid Expansion 0.12 -2.35 -0.49 -2.56 -2.56 -1.37 (3.75) (4.29) (3.83) (1.49) (1.78) (1.62) This table reports the regression 2 $\times$ 2 DiD estimate comparing 978 counties that expanded Medicaid in 2014 with 1,222 counties that did not expand Medicaid by 2019, adjusting for covariates (percent female, percent white, percent hispanic, the unemployment rate, the poverty rate, and median household income). Columns 1-3 report unweighted regression results, while columns 4-6 weight by county population aged 20-64 in 2013. Columns 1 and 4 report results for expansion states without covariates, columns 2 and 5 adjust for the baseline levels of the covariates in 2013, and columns 3 and 6 control for the time-varying covariate values in 2014 and 2013. Standard errors (in parentheses) are clustered at the county level.
@@
Although only one covariate-adjusted estimate in Table 5 is (marginally) statistically significant, the point estimates differ noticeably. In the unweighted case, adjusting for the 2013 levels of the covariates decreases the estimated effect of Medicaid expansion on short-run mortality rates from a point estimate of roughly 0.12 to -2.35. However, if we include their time-varying values instead, we estimate an effect of -0.49, a large difference. We find a similar result when using weighted regressions; while the coefficient remains constant (-2.56) when using 2013 values of the covariates, it attenuates to -1.37 if we use (4.3).
@@
The first thing that is clear from these representations is that because time-invariant variables drop out of equation (4.3), a TWFE specification can account for differential trends related to baseline covariate levels only if they enter as interactions with the post-treatment dummy as in equation (4.4). The exact regression specification, therefore, determines the implied conditional parallel trends assumption. Controlling for annual poverty rates really means controlling for poverty changes, and areas that are poor are not the same as areas that are becoming poor.
@@
Suppose we have decided on which variables to include in a conditional parallel trends assumption and whether to measure them in levels or changes. If Assumptions CPT and SO hold with respect to this set of covariates, does $\beta_{treat}$ recover the $ATT(2)$? In the DiD context, Caetano2024 tackle exactly this question. They show that $\beta_{treat}$ equals a weighted average of conditional average treatment effects, defined as $ATT_{x_{k}}(2)\equiv\mathop{}\!\mathbb{E}_{\omega}\big{[}Y_{i,t=2}(1)-Y_{i,t=2}(0)|D_{i}=1,X_{i}=x_{k}\big{]}$, with weights that may not be convex, plus three bias terms reflecting misspecification either in the set of control variables or the fact that they are included linearly. These conclusions relate to recent findings about the properties of regression estimators in the presence of heterogeneity in other contexts, including instrumental variables (Mogstad_Torgovitsky_2024), cross-sectional designs (Angrist_1998_ECMA; Aronow_Samii_2015; Sloczyski2022; Goldsmith-Pinkham_Hull_Kolesar_2024_AER), and panel data (Goodman2021; deChaisemartin2020; Sun2021; Poirier_Sloczyski_2024).
The weighting results suggest that even when the covariates are correctly selected, measured, and added with the correct functional form, $\beta_{treat}$ could be negative even when $ATT_{X_{i}}(2)$ is positive for all values of covariates. Short of this extreme sign-reversal case, $\beta_{treat}$’s weighting scheme generally does not yield the $ATT(2)$ target parameter and instead puts too much weight on $ATT_{X_{i}}(2)$ for $X_{i}$’s that are relatively uncommon among the treated group relative to the untreated group and puts too little weight on $ATT_{X_{i}}(2)$ for $X_{i}$’s that are relatively common among the treated group relative to the untreated group (Sloczyski2022; Caetano2024).
@@
In other words, (4.3) implicitly rules out that treatment effects can vary across covariate-strata, which makes the weighting issues identified by Caetano2024 irrelevant to the interpretation of $\beta_{treat}$. Research on the Medicaid expansion using data on mortality rates by income, however, shows clear evidence of heterogeneous effects (miller2021medicaid; wyse2024).